し.folloir _Atalisti.OM. ancillang: (s) Te X+-X1.5 Caution:Vefurclain acl cass
Answer #1

We know that for given two normal random variables X and Y, independent of each other s.t.

X \sim N(\mu_X,\sigma^2_X) \ \ \ ; Y \sim N(\mu_Y,\sigma^2_Y) \\ \Rightarrow X \pm Y \sim N(\mu_X \pm \mu_Y , \sigma^2_X + \sigma^2_Y)

(i)

Since, X4 and X1 are independent of each other

\\ \Rightarrow T = X_4-X_1 \sim N(\theta - \theta , 2 + 2) \\ \Rightarrow T \sim N(0,4)

Clearly, the distribution of T is independent of \theta .

Thus, T is ancillary.

(ii)
X_1 \sim N(\theta,2) \\ \Rightarrow 2X_1 \sim N(2\theta,2^2*2 = 8)

Now, since, X2, X3 and 2X1 are independent of each other

T = X_2 + X_3 - 2X_1 \sim N(\theta + \theta - 2\theta, 2 + 2 + 8) \\ \Rightarrow T \sim N(0,12)

Clearly, the distribution of T is independent of \theta .

Thus, T is ancillary.

(iii)

X_4 \sim N(\theta,2) \\ \Rightarrow 2X_4 \sim N(2\theta,2^2*2 = 8)

Now, since, X1, X2, X3 and 2X4 are independent of each other

T = X_1+ X_2 + X_3 + 2X_4 \sim N(\theta + \theta + \theta + 2\theta, 2+2+2+8) \\ \Rightarrow T \sim N(5\theta, 14)

Clearly, the distribution of T is not independent of \theta .

Thus, T is not ancillary.

For any queries, feel free to comment and ask.

If the solution was helpful to you, don't forget to upvote it by clicking on the 'thumbs up' button.